OpenStudy (anonymous):

an unknown number is decreased by five and increased by two . the result is -10

OpenStudy (anonymous):

work backwards. start with -10. If the last step was increase by 2, then subtract. your answer is -7

OpenStudy (calculusxy):

x - 5 + 2 = -10

OpenStudy (anonymous):

x-5+2=-10 +5 -2 +5 -2 x = ?

OpenStudy (anonymous):

so i add 5 both sides ?

OpenStudy (anonymous):

Add 5 to both sides, and subtract 2 from both sides

OpenStudy (anonymous):

an unknown number is multiplied by 3 and divided by -4 . the result in -6

OpenStudy (anonymous):

(3x/-4) = -6 multiply both sides by -4 and solve

OpenStudy (anonymous):